OpenStudy (anonymous):

When an altitude is drawn to the hypotenuse of a right triangle, the lengths of the segments of the hypotenuse are 16 and 25. What is the length of the altitude? 20 25 41 50

OpenStudy (accessdenied):

 We can consider these two triangles as similar triangles and how the ratios of the sides are equal: \[ \frac{a}{16} = \frac{25}{a} \] Here, we just have to solve for the positive a.

OpenStudy (anonymous):

So I crossed multiplied and got 2a = 400. That would then equal 200. But that doesn't seems right?

OpenStudy (accessdenied):

Well, you're right to cross multiply. However, the problem is that: a*a =/= 2a 2a = a + a. a^2 = a*a (a to the power of 2)

OpenStudy (accessdenied):

So, rather than "2a = 400," we have "a^2 = 400" and then we take the positive square root. of both sides to cancel the exponent. :)

OpenStudy (anonymous):

So the answer is 20?

OpenStudy (accessdenied):

Yes. :)